# Rethinking On-policy Optimization for Query Augmentation

### v1

@@ 4.1 Methodology
We begin with a rigorous replication of DeepRetrieval’s experiments. Our setup mirrors the original work, including both sparse and dense retrieval models, where the retriever and corpus together form the “environment” in RL terminology. For sparse retrieval, we use BM25 Robertson et al. (1995), and for dense retrieval, we use task-appropriate retrievers, as detailed in Section˜4.2.
For the prompt-based method (SPQE), we instruct an instruction-following LLM to generate a hypothetical document $d^{H}$ addressing the query $q$, and use the concatenated $(q,d^{H})$ as the augmented query. This zero-shot approach differs from Query2Doc Wang et al. (2023), which relies on few-shot prompting. Compared to the similar zero-shot method HyDE Gao et al. (2023), SPQE skips the followup steps of embedding multiple pseudo documents and aggregating their representations, and is designed to function as a simple, strong baseline for zero-shot query augmentation.
@@ 4.2 Experimental Setup
Datasets, Retrieval and Evaluations. We follow the experimental setup of Jiang et al. (2025) to include evidence-seeking retrieval and ad hoc retrieval datasets. For evidence-seeking retrieval, we use Wikipedia-18333The 2018-12-20 snapshot of English Wikipedia, chunked into non-overlapping 100-word passages Ma et al. (2021) as the retrieval corpus, and include three query collections: NQ Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017) and SQuAD Rajpurkar et al. (2016). For ad hoc retrieval, we use following datasets from BEIR Thakur et al. (2021): FEVER Thorne et al. (2018), HotpotQA Yang et al. (2018), NFCorpus Boteva et al. (2016) and SciFact Wadden et al. (2020), as these datasets include train set for on-policy optimization. We also include evaluation on MS MARCO’s dev set with shallow annotation depth Bajaj et al. (2016), alongside DL19 Voorhees et al. (2020) and DL20 Craswell et al. (2025) with in-depth annotations. Refer dataset statistics to Appendix Appx.˜A.
@@
Dataset Base +3B RL +7B RL SPQE Base +3B RL +7B RL SPQE
@@
For evidence-seeking retrieval, we follow Ma et al. (2021) to measure Hit@20 (answer span hits). We use BM25 as the sparse retriever and E5-base-v2 Wang et al. (2022) as the dense retriever.
For ad hoc retrieval, we use NDCG@10 as the evaluation metric following Thakur et al. (2021). For retrieval, we use their respective corpus, and use BM25 for sparse retrieval and Contriever-msmarco Izacard et al. (2022) as the dense retriever. Note that Jiang et al. (2025) use different dense retrievers for different ad hoc retrieval datasets, but we control the dense retriever model.
Backbone Models. We use Qwen2.5-{3B,7B}-Instruct Qwen (2024) as our initial policy. For implementation of SPQE, we use different LLMs. Our main experimental result is based on proprietary GPT-4o-mini Menick et al. (2024) and we also report results with openweighted GPT-OSS-120B Agarwal et al. (2025) and Qwen3-32B-Instruct Yang et al. (2025).
Implementation Details. We use the original codebase from Jiang et al. (2025) and made several modifications to improve training throughput. Our retrieval environment is based on Pyserini Lin et al. (2021) and Faiss Johnson et al. (2019). We use veRL 0.1 Sheng et al. (2025) as the RL framework and PPO Schulman et al. (2017) as the specific on-policy RL algorithm. We did not include group-based on-policy RL algorithms such as GRPO Shao et al. (2024); Guo et al. (2025), as we note a major bottleneck of our RL formulation lies in the agent-environment interaction part — in our case the retrieval step. Using group-based algorithm increases the retrieval calls by the factor of group size, significantly slowing down the overall optimization process. We leave the investigation of other on-policy algorithms to future work. Without otherwise specified, we use learning rate 1e-6, global train batch size 128, actor and critic mini batch size 32, rollout temperature 0.6. All experiments are performed on 8x or 16x Nvidia A100 GPUs, each with 40GB VRAM.
@@
Dataset Base +3B RL +7B RL SPQE Base +3B RL +7B RL SPQE Base +3B RL +7B RL SPQE
@@
Dense E5-base-v2 Retriever
@@
4.3 Result and Analysis
We show the results of evidence-seeking retrieval and ad hoc retrieval in Table˜1. The data demonstrates that both on-policy optimization and simple prompt-based expansion (SPQE) significantly improve retrieval performance over not using any query augmentation (the “Base” columns).
First, the results confirm the value of using on-policy optimization. For ad hoc retrieval with a dense retriever, the +7B RL model achieves the highest average score of 57.5, a substantial improvement over the base retriever’s score of 51.6. This gain is particularly noticeable on datasets like FEVER, where the score increases from 75.8 to 85.5. Similarly, with a sparse BM25 retriever on evidence-based datasets, the +7B RL model improves the average score from 70.2 to 75.6. These results show that directly training a model to generate queries that maximize retrieval scores is an effective strategy.
Furthermore, Table˜1 shows that our simpler prompt-based method, SPQE, is also highly effective and competitive. For sparse retrieval, SPQE consistently delivers the best performance, achieving the highest average scores for both evidence-based (78.2) and ad hoc (51.6) datasets. For dense retrieval, its performance is very close to the more complex RL method, even outperforming it on several individual ad hoc datasets like DL19 (70.1) and DL20 (69.3). Additionally, we note that SPQE is flexible about LLMs’ choices, as Qwen-3-32B-Instruct and GPT-OSS-120B also achieve strong performance (details in Appendix Appx.˜B). We defer a more in-depth discussion about SPQE vs. on-policy optimization to Section˜4.5.
4.4 Extending to Tool Retrieval
@@
Retrieving useful tools from tool sets for LLM agents is becoming increasingly critical in real-world applications. However, the main challenge is that candidate tools are usually large-scale and many of them are similar in functionality Patil et al. (2024); Qin et al. (2024); Qu et al. (2024); Shi et al. (2025). How query augmentation might enhance tool retrieval efficacy is rather underexplored due to the lack of large-scale training and evaluation datasets. We address the gap by studying prompting-based and RL-based query augmentation in the context of tool retrieval.
@@
Datasets. We use a recently released tool retrieval dataset ToolRet Shi et al. (2025), which is constructed from existing tool learning datasets. The train set contains over 200k instances, which allows us to study on-policy RL training at scale.
@@
Training and Evaluation. We construct the trainset and train corpus from the released trainset. Note that here the train corpus (i.e., environment) is different from the corpus we run evaluation on, due to the mismatch of candidate tools. We evaluate on three official splits: ToolRet-Web, ToolRet-Code, ToolRet-Customized. We train the same base policy as Section˜4.2: Qwen2.5-{3B, 7B}-Instruct and adopt BM25 and E5-base-v2 as the retrievers. We report performance of SPQE as a simple prompting-based method compared to on-policy optimization. For evaluation, we focus on classical retrieval metrics NDCG@10 and Recall@10. We further include Completeness@10. The Completeness metric was proposed by Qu et al. (2024) to measure whether all the required tools have been included in the ranklist. For reward design, we directly optimize against the Completeness@10 metric. We leave further investigation of reward design to future works.
@@
Results and Analysis. We show the results of tool retrieval in Table˜2. Similar to our prior results, we observe that query augmentation methods can substantially improve performance over the baseline. On-policy optimization shows some benefits, particularly for the sparse retriever, but also reveals a notable disparity in its effectiveness between sparse and dense models.
@@
For Completeness@10, the metric we optimized for, the +3B RL model improves the average score for the sparse BM25 retriever from a base of 29.4 to 32.3. This demonstrates that RL can successfully steer the model to improve a target metric. However, for the dense retriever, the same RL training shows no benefit, with average Completeness@10 scores (29.9 for +3B RL, 29.4 for +7B RL) failing to surpass the base score of 29.6. This disparity between sparse and dense retriever is critical. While RL provides a modest lift for BM25 on both NDCG@10 (from 30.0 to 32.0) and Completeness@10, it consistently degrades performance for the dense Contriever across all metrics. For instance, the +7B RL model lowers the average NDCG@10 from 29.9 to 29.0. This observation suggests that the queries generated by the RL policy, while optimized for the Completeness reward, may be producing keyword-focused outputs that help term-matching systems like BM25 but result in less effective embeddings for semantic-focused dense retrievers.
@@
In contrast, the simple prompting method SPQE proves to be the most effective and robust method across the board. It achieves the highest average scores for both NDCG@10 (33.4 for sparse, 34.5 for dense) and Completeness@10 (35.1 for sparse, 34.8 for dense). Furthermore, looking at Recall@10, HyDE, another generative approach, consistently outperforms all other methods. This indicates that for the task of tool retrieval, generating richer, document-like queries via simple prompting is a more reliable strategy for improving performance on both sparse and dense systems than the RL approach tested here.
@@ 4.5 Discussions
In contrast, prompt-based expansion (e.g., SPQE) is a robust and resource-efficient baseline that performs well across nearly all settings. We hypothesize that its effectiveness stems from three factors. First, it leverages the LLM’s internal knowledge to expand queries with relevant terms, mitigating the vocabulary mismatch problem for sparse retrievers like BM25. Second, it transforms asymmetric query-document retrieval into a symmetric document-document task, which better aligns with the training objectives of many dense retrievers and boosts their performance. Third, unlike RL methods that rewrite the original query, prompt-based expansion produces a pseudo-document, offering richer context for the retriever to match against. We hypothesize this additional signal makes pseudo-documents especially effective on existing IR benchmarks.
@@
Dataset Base +3B +7B SPQE OPQE-3B OPQE-7B Base +3B +7B SPQE OPQE-3B OPQE-7B

### v2

@@ 4.1 Methodology
We begin with a rigorous replication of DeepRetrieval’s experiments. Our setup mirrors the original work, including both sparse and dense retrieval models, where the retriever and corpus together form the “environment” in RL terminology. For sparse retrieval, we use BM25 (Robertson et al., 1995), and for dense retrieval, we use task-appropriate retrievers, as detailed in Section˜4.2. In tables, we denote the DeepRetrieval baselines as DR-3B and DR-7B based on the policy size.
Replication Checklist.
To ensure faithful replication, we follow DeepRetrieval’s overall formulation: (i) the policy generates intermediate reasoning followed by a final augmented query in a fixed format (Appendix Appx.˜E); (ii) the retrieval reward is computed from the final query against a fixed retriever–corpus environment; (iii) the overall reward includes a format constraint and a retrieval component (Appendix Appx.˜C); and (iv) we optimize with PPO under a KL-regularized objective (Algorithm˜1). We only modify engineering details to improve throughput (e.g., environment implementation), while keeping the task definition, reward, and evaluation metrics aligned with the original setup.
For the prompt-based method (SPQE), we instruct an instruction-following LLM to generate a hypothetical document $d^{H}$ addressing the query $q$, and use the concatenated $(q,d^{H})$ as the augmented query. This zero-shot approach differs from Query2Doc (Wang et al., 2023), which relies on few-shot prompting. Compared to the similar zero-shot method HyDE Gao et al. (2023), SPQE skips the follow-up steps of embedding multiple pseudo-documents and aggregating their representations, and is designed to function as a simple, strong baseline for zero-shot query augmentation.
@@ 4.2 Experimental Setup
Datasets, Retrieval and Evaluations.
We follow the experimental setup of Jiang et al. (2025) to include evidence-seeking retrieval and ad hoc retrieval datasets. For evidence-seeking retrieval, we use Wikipedia-18333The 2018-12-20 snapshot of English Wikipedia, chunked into non-overlapping 100-word passages (Ma et al., 2021) as the retrieval corpus, and include three query collections: NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017) and SQuAD (Rajpurkar et al., 2016). For ad hoc retrieval, we use following datasets from BEIR (Thakur et al., 2021): FEVER (Thorne et al., 2018), HotpotQA (Yang et al., 2018), NFCorpus (Boteva et al., 2016) and SciFact (Wadden et al., 2020), as these datasets include train set for on-policy optimization. We also include evaluation on MS MARCO’s dev set with shallow annotation depth (Bajaj et al., 2016), alongside DL19 (Voorhees et al., 2020) and DL20 (Craswell et al., 2025) with in-depth annotations. Refer dataset statistics to Appendix Appx.˜A.
For evidence-seeking retrieval, we follow Ma et al. (2021) to measure Hit@20 (answer span hits). We use BM25 as the sparse retriever and E5-base-v2 (Wang et al., 2022) as the dense retriever.
For ad hoc retrieval, we use NDCG@10 as the evaluation metric following Thakur et al. (2021). For retrieval, we use their respective corpus, and use BM25 for sparse retrieval and Contriever-msmarco (Izacard et al., 2022) as the dense retriever. Note that Jiang et al. (2025) use different dense retrievers for different ad hoc retrieval datasets, but we control the dense retriever model.
Backbone Models.
We use Qwen2.5-{3B,7B}-Instruct (Qwen, 2024) as our initial policy. For implementation of SPQE, we use different LLMs. Our main experimental result is based on proprietary GPT-4o-mini (Menick et al., 2024) and we also report results with open-weight GPT-OSS-120B (Agarwal et al., 2025) and Qwen3-32B-Instruct (Yang et al., 2025). Importantly, we intentionally allocate more inference-time compute to prompting by using larger LLMs, so that the overall compute budget is closer to that of on-policy RL training (which uses smaller base policies but incurs substantial training-time compute). As a result, our SPQE baselines are not backbone-matched to the DR-3B/7B policies; in preliminary experiments, SPQE with the same Qwen2.5-{3B,7B}-Instruct backbones was not as competitive, as they are not directly trained for the query rewriting tasks. Concretely, our PPO training runs require multi-GPU A100 setups (e.g., 7-14x A100 40GB), whereas running a 32B instruction model for SPQE inference can be done on a single A100 40GB GPU (e.g., using vLLM). Note that SPQE incurs per-query inference cost, while RL incurs an up-front multi-GPU training cost that is amortized over the evaluation set. Therefore, their wallclock costs are not directly comparable. In practice, when per-dataset RL training is infeasible (e.g., due to compute or frequent domain shifts), a zero-shot prompting baseline like SPQE can serve as a strong default.
To reduce the possibility that SPQE performance is driven solely by access to proprietary LLMs, we additionally evaluate SPQE with strong open-weight instruction models (Qwen3-32B-Instruct and GPT-OSS-120B). As shown in Table˜2, these open-weight models achieve performance that is generally comparable to GPT-4o-mini across evidence-seeking, ad hoc, and tool retrieval settings (full per-dataset results in Appendix Appx.˜B). Notably, SPQE with an open-weight 32B model can be served on a single A100 40GB GPU (e.g., using vLLM).
Implementation Details.
We use the original codebase from Jiang et al. (2025) and made several modifications to improve training throughput. Our retrieval environment is based on Pyserini (Lin et al., 2021) and Faiss (Johnson et al., 2019). We use veRL 0.1 (Sheng et al., 2025) as the RL framework and PPO (Schulman et al., 2017) as the specific on-policy RL algorithm, matching DeepRetrieval’s PPO-based training setup. We did not include group-based on-policy RL algorithms such as GRPO (Shao et al., 2024; Guo et al., 2025); while such alternatives may yield similar performance in single-turn retrieval environments with verifiable rewards, we do not evaluate them here and leave a systematic comparison of on-policy algorithms to future work. Without otherwise specified, we use learning rate 1e-6, global train batch size 128, actor and critic mini batch size 32, rollout temperature 0.6. All experiments are performed on 8x or 16x Nvidia A100 GPUs, each with 40GB VRAM.
@@
Dataset Base DR-3B DR-7B SPQE Base DR-3B DR-7B SPQE
@@
Table 2: Prompting (SPQE) with proprietary vs. open-weight LLMs, compared against RL-trained policies (averaged over datasets). Evidence-seeking uses Hit@20, ad hoc retrieval uses NDCG@10, and tool retrieval uses Recall@10. Full per-dataset results for SPQE are provided in Appendix Appx.˜B.
Task family Retriever Base GPT-4o-mini Qwen3-32B-Instruct GPT-OSS-120B DR-3B DR-7B
Evidence-seeking BM25 70.2 78.2 76.5 80.0 75.0 75.6
Evidence-seeking E5-base-v2 79.9 79.7 78.5 79.7 79.6 79.8
Ad hoc (BEIR) BM25 46.6 51.6 50.8 51.3 48.4 49.0
Ad hoc (BEIR) Contriever 56.0 56.6 56.2 51.8 57.3 57.5
Tool retrieval BM25 39.9 44.4 44.1 44.3 41.4 39.8
Tool retrieval Contriever 38.8 44.8 45.4 44.3 38.6 37.8
Overall
Borda score Sparse 4.0 17.0 12.0 16.0 7.0 7.0
Borda score Dense 11.0 12.5 10.0 8.5 9.0 12.0
Borda score All 15.0 29.5 22.0 24.5 16.0 19.0
Borda rank All 6 1 3 2 5 4
4.3 Result and Analysis
We show the results of evidence-seeking retrieval and ad hoc retrieval in Table˜1. Overall, both on-policy optimization and simple prompt-based expansion (SPQE) improve retrieval performance over not using any query augmentation (the “Base” columns), but their benefits differ across tasks and retrievers.
On-policy optimization can be effective, especially when training smaller policies. For ad hoc retrieval with a dense retriever, the DR-7B (DeepRetrieval) model achieves the highest average score of 57.5, improving over the base retriever’s score of 51.6; this is particularly noticeable on datasets like FEVER, where the score increases from 75.8 to 85.5. Similarly, with a sparse BM25 retriever on evidence-based datasets, the DR-7B model improves the average score from 70.2 to 75.6.
Furthermore, Table˜1 shows that our simpler prompt-based method, SPQE, is also highly effective and competitive. For sparse retrieval, SPQE consistently delivers the best performance, achieving the highest average scores for both evidence-based (78.2) and ad hoc (51.6) datasets.
To provide a consolidated view across sparse vs. dense settings (and to disentangle backbone effects for prompting), we summarize prompting with proprietary and open-weight LLMs alongside the DeepRetrieval baselines in Table˜2. Across all six task–retriever combinations, prompting achieves the best overall Borda rank (rank 1), while the DeepRetrieval policies rank behind the prompting variants (rank 4 for DR-7B and rank 5 for DR-3B). This aggregate result reinforces the main takeaway from Table˜1: under our compute-aware setup (larger inference-time models for prompting vs. smaller RL-trained policies with substantial training cost), simple pseudo-document prompting is a very strong baseline and achieves the best overall average ranking in our experiments.
For dense retrieval, prompting performance is also close to (and sometimes surpasses) the more complex RL method, while RL can still be beneficial on specific datasets (e.g., FEVER and FiQA) when adapted to a particular retriever and domain. We defer a more in-depth discussion about SPQE vs. on-policy optimization to Section˜4.5.
4.4 Extending to Tool Retrieval
@@
Dataset Base DR-3B DR-7B SPQE Base DR-3B DR-7B SPQE Base DR-3B DR-7B SPQE
@@
Dense Contriever-msmarco Retriever
@@
Retrieving useful tools from tool sets for LLM agents is becoming increasingly critical in real-world applications. However, the main challenge is that candidate tools are usually large-scale and many of them are similar in functionality (Patil et al., 2024; Qin et al., 2024; Qu et al., 2024; Shi et al., 2025). How query augmentation might enhance tool retrieval efficacy is rather underexplored due to the lack of large-scale training and evaluation datasets. We address the gap by studying prompting-based and RL-based query augmentation in the context of tool retrieval.
Datasets.
@@
We use a recently released tool retrieval dataset ToolRet (Shi et al., 2025), which is constructed from existing tool learning datasets. The train set contains over 200k instances, which allows us to study on-policy RL training at scale.
Training and Evaluation.
@@
We construct the trainset and train corpus from the released trainset. Note that here the train corpus (i.e., environment) is different from the corpus we run evaluation on, due to the mismatch of candidate tools. We evaluate on three official splits: ToolRet-Web, ToolRet-Code, ToolRet-Customized. We train the same base policy as Section˜4.2: Qwen2.5-{3B, 7B}-Instruct and adopt BM25 and E5-base-v2 as the retrievers. We report performance of SPQE as a simple prompting-based method compared to on-policy optimization. For evaluation, we focus on classical retrieval metrics NDCG@10 and Recall@10. We further include Completeness@10. The Completeness metric was proposed by Qu et al. (2024) to measure whether all the required tools have been included in the ranklist. For reward design, we directly optimize against the Completeness@10 metric. We leave further investigation of fine-grained reward design to future works.
Results and Analysis.
@@
We show the results of tool retrieval in Table˜3. Similar to our prior results, we observe that query augmentation methods can substantially improve performance over the baseline. On-policy optimization shows some benefits, particularly for the sparse retriever, but also reveals a notable disparity in its effectiveness between sparse and dense models.
@@
For Completeness@10, the metric we optimized for, the DR-3B (DeepRetrieval) model improves the average score for the sparse BM25 retriever from a base of 29.4 to 32.3. This demonstrates that RL can successfully steer the model to improve a target metric. However, for the dense retriever, the same RL training shows no benefit, with average Completeness@10 scores (29.9 for DR-3B, 29.4 for DR-7B) failing to surpass the base score of 29.6. This disparity between sparse and dense retriever is critical. While DeepRetrieval provides a modest lift for BM25 on both NDCG@10 (from 30.0 to 32.0) and Completeness@10, it consistently degrades performance for the dense Contriever across all metrics. For instance, the DR-7B model lowers the average NDCG@10 from 29.9 to 29.0. This observation suggests that the queries generated by the RL policy, while optimized for the Completeness reward, may be producing keyword-focused outputs that help term-matching systems like BM25 but result in less effective embeddings for semantic-focused dense retrievers. A more direct diagnosis would require analyzing model outputs (e.g., length/verbosity, term distribution shifts, and query–tool coverage); we do not include these statistics in this paper and leave them to future work.
@@
In contrast, the simple prompting method SPQE is consistently strong across both sparse and dense tool retrieval settings. It achieves the highest average scores for both NDCG@10 (33.4 for sparse, 34.5 for dense) and Completeness@10 (35.1 for sparse, 34.8 for dense). Furthermore, looking at Recall@10, HyDE, another generative approach, consistently outperforms all other methods. This indicates that for the task of tool retrieval, generating richer, document-like queries via simple prompting can be a reliable strategy for improving performance on both sparse and dense systems relative to the specific RL formulation tested here.
@@ 4.5 Discussions
In contrast, prompt-based expansion (e.g., SPQE) is a robust and resource-efficient baseline that performs well across nearly all settings. We hypothesize that its effectiveness stems from three factors. First, it leverages the LLM’s internal knowledge to expand queries with relevant terms, mitigating the vocabulary mismatch problem for sparse retrievers like BM25; this is consistent with the strong zero-shot BM25 gains in Table˜1 and the higher initial rewards when training OPQE (see Fig.˜1). Second, it transforms asymmetric query-document retrieval into a symmetric document-document task, which better aligns with the training objectives of many dense retrievers and boosts their performance. Third, unlike RL methods that rewrite the original query, prompt-based expansion produces a pseudo-document, offering richer context for the retriever to match against. We hypothesize this additional signal makes pseudo-documents especially effective on existing IR benchmarks.
@@
Dataset Base DR-3B DR-7B SPQE OPQE-3B OPQE-7B Base DR-3B DR-7B SPQE OPQE-3B OPQE-7B
@@
5 On-policy Pseudo-document Query Expansion
5.1 Method
We make a simple modification to the DeepRetrieval style on-policy optimization RL process for direct query rewriting. Instead of instructing the policy model to rewrite the query $q\rightarrow q^{\prime}$, we instruct the model to write a hypothetical document $d^{H}$, then directly output the concatenated $(q,d^{H})$ (prompt template in Appendix Appx.˜E). This reformulation inherently changes the output structure and typical length distribution of the augmented query compared to direct query rewriting. Then this concatenated string is used to retrieve relevant documents from the environment, similar to SPQE, and the reward is computed by the retrieval metrics like NDCG. In this way, we merge the SPQE into the on-policy optimization process. We refer to this method as On-policy Pseudo-document Query Expansion (OPQE).
5.2 Experiment, Result and Analysis
Experimental Setup.
We experiment on evidence-seeking retrieval and ad hoc retrieval datasets. Our implementation is the same as Section˜4.2 except for the prompt template to instruct the policy model to write a pseudo-document.
Main Results.
We report the results of our hybrid approach in Table˜4. The key finding is that our proposed method, OPQE, successfully integrates the strengths of both prompting and on-policy optimization, achieving the best overall performance, especially in the dense retrieval setting.